Jefferson Lab HALL A Logbook for February 01 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 57204 02/28/01 22:59 Higinbotham Harp Scan and OTR
  • 57200 02/28/01 22:14 Higinbotham BeO Check
  • 57193 02/28/01 19:44 Lingyan E_gamma for coin15_LD2
  • 57191 02/28/01 19:07 vasulk Reboot IOC
  • 57190 02/28/01 19:04 Higinbotham Magnet Screen
  • 57185 02/28/01 16:55 jpchen flow plot from cryo group
  • 57184 02/28/01 16:47 jpchen Guideline for change LH2/LD2 fan frequency
  • 57181 02/28/01 16:37 jpchen LH2 fan heat power test
  • 57180 02/28/01 16:31 jpchen Target operators, please adjust coolant flow (J-T valve) slowly
  • 57178 02/28/01 16:01 LeRose Shift Summary
  • 57173 02/28/01 14:30 Lingyan LH2 fan problem
  • 57169 02/28/01 13:15 LeRose Adjusted Collimators
  • 57168 02/28/01 13:15 Ole Hansen Changed X Windows fonts on hacsbc2
  • 57166 02/28/01 13:07 Ole Hansen Radiator is indeed out
  • 57162 02/28/01 11:38 Lingyan harp scan
  • 57161 02/28/01 11:30 LeRose Magnet Screen
  • 57160 02/28/01 11:26 Saha Arc Integral Measurement
  • 57159 02/28/01 11:26 LeRose BeO Spot
  • 57158 02/28/01 09:34 atrig Trigger-Download
  • 57157 02/28/01 08:33 Dutta shift summary
  • 57132 02/28/01 02:30 Adaq Egamma spectrum for LD2 (coin 13)
  • 57130 02/28/01 02:27 Adaq Q3 tripped off 3rd time tonight
  • 57122 02/28/01 00:05 Higinbotham Shift Summary
  • Tuesday

  • 57121 02/27/01 23:58 Higinbotham Q3 tripped off again
  • 57120 02/27/01 23:40 Higinbotham Q3 coming back online
  • 57119 02/27/01 23:30 Higinbotham Right Q3 Tripped Off
  • 57113 02/27/01 22:37 Higinbotham Spot Check
  • 57107 02/27/01 22:01 reitz BeO postion check
  • 57106 02/27/01 21:58 Higinbotham Harp Scan
  • 57103 02/27/01 20:11 atrig Trigger-Download
  • 57100 02/27/01 18:10 jpchen set target to LD2 running condition
  • 57099 02/27/01 18:07 jpchen LH2/LD2 operation: guideline for target operators
  • 57098 02/27/01 17:21 Lingyan scintillator efficiency for coin13_he4
  • 57097 02/27/01 17:13 jpchen/mccor Target cooldown completed
  • 57096 02/27/01 16:29 jpchen He4 target pressure transducer calibrated
  • 57095 02/27/01 16:28 jpchen BeO target in position
  • 57094 02/27/01 16:27 jpchen Cryotarget switch from He4 to LH2/LD2 status
  • 57092 02/27/01 16:11 jpchen/mccor Target motion during warm-up and cooldown.
  • 57091 02/27/01 15:52 atrig Trigger-Download
  • 57090 02/27/01 12:39 X. Zheng Gas Cherenkov Calibration
  • 57089 02/27/01 11:50 Jack Gas sensors.
  • 57087 02/27/01 07:55 ddutta Shift summary
  • 57047 02/27/01 00:41 Lingyan Zhu swing shift summary
  • Monday

  • 57043 02/26/01 23:53 Reitz bpm/harp calibration pedestals
  • 57042 02/26/01 23:49 Lingyan target IOC reboot
  • 57037 02/26/01 22:47 Lingyan E_gamma for He4 coin13
  • 57021 02/26/01 18:45 Lingyan beam positon on BeO
  • 57020 02/26/01 18:18 Lingyan r-arm dipole ramp curve
  • 57019 02/26/01 18:17 Jack Right Arm flammable gas sensor.
  • 57018 02/26/01 18:06 Lingyan One L-arm bad cable replaced
  • 57014 02/26/01 15:40 atrig Trigger-Download
  • 57013 02/26/01 15:06 weinstein shift summary
  • 57007 02/26/01 14:01 R. Michaels how good is DAQ-busy deadtime
  • 57005 02/26/01 13:26 Calarco Magnet ramp down
  • 57004 02/26/01 12:19 R. Michaels fixed a few scaler problems
  • 56992 02/26/01 07:55 ddutta Shift Summary
  • 56983 02/26/01 05:35 feng Coda crash
  • 56978 02/26/01 04:17 Adaq radiator is out
  • 56969 02/26/01 00:28 Lingyan Swing shift summary
  • Sunday

  • 56965 02/25/01 22:13 Branko BCM coefficients
  • 56959 02/25/01 20:14 Lingyan Beam position on BeO
  • 56957 02/25/01 18:35 Calarco Close call!!!
  • 56956 02/25/01 18:18 Calarco Right dipole (continued)
  • 56955 02/25/01 18:12 Calarco Magnet strip chart
  • 56954 02/25/01 18:09 Calarco Procedure for right dipole
  • 56953 02/25/01 17:46 Calarco How to set the right dipole
  • 56950 02/25/01 14:53 atrig Trigger-Download
  • 56946 02/25/01 13:02 Calarco/Dohr Magnet Quench Right D
  • 56945 02/25/01 11:42 jpchen Completed He4 target cool-down
  • 56944 02/25/01 10:37 Calarco Hall Status
  • 56943 02/25/01 08:49 Calarco Magnets
  • 56938 02/25/01 07:48 Feng Shift summary
  • 56937 02/25/01 07:09 jpchen Target lift power reset
  • 56936 02/25/01 07:05 jpchen cryotarget cooldown
  • Saturday

  • 56932 02/24/01 19:22 atrig Trigger-Download
  • 56929 02/24/01 18:13 Lingyan DAQ not completely back
  • 56922 02/24/01 15:05 Calarco Shift summary
  • 56921 02/24/01 12:44 Calarco Power Failure
  • 56920 02/24/01 12:36 Adaq power outage
  • 56909 02/24/01 09:24 dohrmann Left Arm D1 NMR was locked
  • 56907 02/24/01 07:51 R. Michaels owl shift summary
  • 56906 02/24/01 07:11 R. Michaels magnet status, no detector access
  • 56905 02/24/01 04:45 R. Michaels R-arm dipole quench, He level low
  • 56904 02/24/01 04:14 atrig Trigger-Download
  • 56903 02/24/01 03:49 R. Michaels switch to kine 14
  • 56900 02/24/01 02:18 Lingyan prescale and deadtime error from run 2159 to run 2169
  • 56897 02/24/01 02:03 R. Michaels re: odd mult. peaks in TDCs
  • 56888 02/24/01 00:00 dohrmann Shift summary
  • Friday

  • 56883 02/23/01 22:06 Adaq rebooted rocs 1,2,14, restarted coda
  • 56879 02/23/01 21:20 adaq some beamtrips
  • 56867 02/23/01 17:49 adaq beam back at 17:25
  • 56865 02/23/01 17:14 adaq no beam till at least 17:30
  • 56864 02/23/01 17:02 Adaq no beam due to MCC studies
  • 56857 02/23/01 14:45 reimer trigger for 2154-56
  • 56855 02/23/01 14:40 atrig Trigger-Download
  • 56854 02/23/01 14:39 atrig Trigger-Download
  • 56848 02/23/01 13:53 gilman pulsed beam again
  • 56846 02/23/01 13:05 reimer BeO beam position
  • 56845 02/23/01 12:27 gilman loop1 fan frequency
  • 56843 02/23/01 11:44 gilman cosmic test run, wrong kinematics
  • 56841 02/23/01 11:35 jpchen jpchen
  • 56840 02/23/01 11:28 gilman spectrometers set for 4.3 GeV
  • 56839 02/23/01 11:15 jpchen TO needs watch pressure/temperature and heat load carefully
  • 56838 02/23/01 11:10 jpchen BeO target in position
  • 56837 02/23/01 11:06 jpchen He4 re-cool-down completed
  • 56836 02/23/01 10:05 gilman what is happening now
  • 56834 02/23/01 08:24 Hong Xiang Owl shift6 Summary(2/23)
  • 56833 02/23/01 07:42 Lingyan target vent
  • 56819 02/23/01 03:57 Hong Xiang Difference in He4 and LD2 Spectra?
  • 56815 02/23/01 02:47 atrig Trigger-Download
  • 56802 02/23/01 00:04 Lingyan Zhu Swing shift summary
  • Thursday

  • 56798 02/22/01 23:13 atrig Trigger-Download
  • 56790 02/22/01 20:56 Lingyan Target IOC reboot
  • 56789 02/22/01 20:49 Lingyan E_gamma for He4 coin9
  • 56786 02/22/01 20:13 R. Michaels pulser is double counting
  • 56777 02/22/01 17:43 Kumbartzki BeO spot check/ Harp
  • 56776 02/22/01 17:39 Hong Xiang 2/22/01 Day Shift Summary
  • 56770 02/22/01 16:31 atrig Trigger-Download
  • 56769 02/22/01 16:28 jpchen Need to set to run condition for HE4 before beam come.
  • 56768 02/22/01 16:22 jpchen Target motion during warmup and cooldown
  • 56767 02/22/01 16:13 x. jiang djust pion rejector first layer high voltages again
  • 56766 02/22/01 16:03 X. Jiang left-arm high voltage screen at run 2125
  • 56765 02/22/01 16:03 jpchen cryotarget switch from LH2/LD2 to He4 completed.
  • 56761 02/22/01 09:49 Hong Xiang Spec setting for survey
  • 56759 02/22/01 06:39 Kumbartzki Owl shift summary
  • 56726 02/22/01 00:53 atrig Trigger-Download
  • 56722 02/22/01 00:07 Hong Xiang Shift Summary(swing)
  • Wednesday

  • 56701 02/21/01 21:49 Hong Xiang Egamma Spectra, sing9_rev Setting
  • 56687 02/21/01 20:43 atrig Trigger-Download
  • 56666 02/21/01 16:06 atrig Trigger-Download
  • 56664 02/21/01 16:01 Adaq Day shift summary
  • 56661 02/21/01 15:27 Kumbartzki BeO check
  • 56660 02/21/01 14:33 X. Jiang changed left -arm pion rejector High Voltage settings
  • 56659 02/21/01 14:21 X. Jiang Left -arm high voltage screen after run 2078
  • 56658 02/21/01 14:04 stevensm right dipole locked up
  • 56657 02/21/01 13:52 Jack Right arm gas sensor.
  • 56652 02/21/01 08:02 Hong Xiang Shift Summary
  • 56650 02/21/01 07:58 atrig Trigger-Download
  • 56627 02/21/01 04:17 atrig Trigger-Download
  • 56625 02/21/01 04:09 kramerk left-gas cerenkov plot
  • 56623 02/21/01 04:01 bogdanw HV on Left Arm GC
  • 56613 02/21/01 01:05 Lingyan target density test run summary
  • 56602 02/21/01 00:06 Kumbartzki Swing shift summary
  • Tuesday

  • 56571 02/20/01 20:24 bogdanw Left Arm GC #9
  • 56570 02/20/01 20:22 atrig Trigger-Download
  • 56569 02/20/01 20:18 atrig Trigger-Download
  • 56568 02/20/01 20:17 atrig Trigger-Download
  • 56549 02/20/01 16:02 Olivier Day shift summary
  • 56547 02/20/01 15:51 atrig Trigger-Download
  • 56546 02/20/01 15:47 Saha Arc Integral measurement
  • 56545 02/20/01 15:27 atrig Trigger-Download
  • 56544 02/20/01 15:08 Olivier Ramping up magnets
  • 56543 02/20/01 12:52 Saha Arc measurement in Achromatic Mode
  • 56541 02/20/01 12:18 Saha Harp Scans
  • 56540 02/20/01 12:11 Olivier BeO target check
  • 56521 02/20/01 09:43 Olivier Beam off
  • 56519 02/20/01 09:21 Olivier ROC 14 reset
  • 56517 02/20/01 09:07 Olivier DAQ problem
  • 56505 02/20/01 07:57 Hong Xiang Shift Summary
  • 56493 02/20/01 06:44 atrig Trigger-Download
  • 56486 02/20/01 05:42 Hong Xiang Gas System Alarm went off
  • 56485 02/20/01 04:38 atrig Trigger-Download
  • 56451 02/20/01 00:47 atrig Trigger-Download
  • 56445 02/20/01 00:07 Kumbartzki Swing shift summary
  • Monday

  • 56422 02/19/01 20:28 Lingyan egamma of sing9
  • 56421 02/19/01 20:28 K. McCormick ESR Temperature high
  • 56418 02/19/01 20:26 Lingyan E_gamma plot for Coin7
  • 56407 02/19/01 18:52 Adaq Harp scan, again
  • 56405 02/19/01 18:47 Adaq Harp Scan
  • 56404 02/19/01 16:05 S. Strauch Schift summjary
  • 56402 02/19/01 13:55 Wang Xu The Upgraded Espace
  • 56401 02/19/01 10:24 jpchen Moved target to BeO
  • 56400 02/19/01 09:00 Lingyan E_gamma from LD2 sing9
  • 56399 02/19/01 08:01 Lingyan shift summary
  • 56397 02/19/01 06:42 Lingyan Beam off
  • 56392 02/19/01 05:27 Lingyan FF on before 1964
  • 56391 02/19/01 04:20 atrig Trigger-Download
  • 56380 02/19/01 02:25 Lingyan E_gamma spectrum of coin7
  • 56375 02/19/01 00:14 atrig Trigger-Download
  • Sunday

  • 56372 02/18/01 23:59 Ole Hansen Shift Summary
  • 56371 02/18/01 23:55 Ole Hansen FF off. Beam position drifting.
  • 56358 02/18/01 20:18 H. Xiang E_gamma spectrum for coin8
  • 56352 02/18/01 18:45 Ole Hansen Runs 1951-1952 are junk
  • 56346 02/18/01 17:25 atrig Trigger-Download
  • 56345 02/18/01 16:03 Olivier Day shift summary
  • 56335 02/18/01 14:22 Olivier BeO target
  • 56334 02/18/01 14:19 OIivier CODA in unbuffered mode
  • 56333 02/18/01 14:03 Olivier Start run comments correction
  • 56324 02/18/01 12:30 H. Xiang E_gamma spectrum for coin9
  • 56323 02/18/01 12:09 Olivier Beam dump problem
  • 56322 02/18/01 11:33 Z.-E. Mezian Cryotarget
  • 56315 02/18/01 09:49 Olivier L- Gas Cerenkov HV
  • 56311 02/18/01 08:23 Olivier Optics data
  • 56308 02/18/01 08:04 Lingyan Shift summary
  • 56302 02/18/01 05:08 Lingyan TDC roc2 slot 4 ok after suppress the pedestals
  • 56291 02/18/01 03:52 atrig Trigger-Download
  • 56289 02/18/01 02:50 Lingyan R-arm Q2 fixed
  • 56287 02/18/01 00:12 R. Michaels adaqs2 root partition filled up, ok now
  • Saturday

  • 56286 02/17/01 23:36 Gao R-arm Q2 quenched again!
  • 56285 02/17/01 21:00 Adaq Pulse beam!
  • 56284 02/17/01 19:42 H. Gao Right-arm Q2 quench!
  • 56283 02/17/01 17:56 atrig Trigger-Download
  • 56282 02/17/01 17:51 atrig Trigger-Download
  • 56281 02/17/01 13:46 Olivier radiator motion
  • 56278 02/17/01 11:13 gomez EPICS control screens
  • 56277 02/17/01 03:55 Hong Xiang EPICS HV Control Display Problem
  • 56276 02/17/01 02:29 Lingyan BPM/Harp calibration results
  • 56273 02/17/01 01:11 Lingyan l-arm VDC: missing wires back
  • Friday

  • 56271 02/16/01 15:57 atrig Trigger-Download
  • 56270 02/16/01 15:51 R. Michaels TDC roc 1 slot 21 ok now
  • 56267 02/16/01 14:29 Mitchell/Hol target cool down motion
  • Thursday

  • 56257 02/15/01 20:33 Lingyan PMT and HV change to l-arm gas#7,#9 and s1#1r, r-arm S1#3l
  • 56254 02/15/01 16:58 R. Michaels extra gates
  • 56253 02/15/01 16:36 R. Michaels missing gates, "fixed"
  • 56249 02/15/01 14:33 R. Michaels put T1,T3 into TDC as flag
  • Wednesday

  • 56242 02/14/01 14:19 R. Michaels files split by hand
  • 56240 02/14/01 12:02 atrig Trigger-Download
  • Tuesday

  • 56239 02/13/01 21:43 Lingyan scintillator threhold
  • 56238 02/13/01 21:30 Lingyan left-scintillator calibration
  • 56237 02/13/01 20:29 Lingyan right arm current
  • 56236 02/13/01 20:19 Lingyan l-arm HV setting
  • 56231 02/13/01 13:14 Lingyan left-HV 4.8?
  • 56230 02/13/01 13:09 Lingyan Left arm HV restored and Right arm problem
  • 56229 02/13/01 13:03 Lingyan A2 #25 and #26 pedestal narrowed
  • 56228 02/13/01 12:47 atrig Trigger-Download
  • Wednesday

  • 56224 02/07/01 11:45 R. Michaels trigger restored
  • 56223 02/07/01 11:25 atrig Trigger-Download
  • 56222 02/07/01 11:25 atrig Trigger-Download
  • 56221 02/07/01 11:24 atrig Trigger-Download
  • 56219 02/07/01 10:10 atrig Trigger-Download
  • Tuesday

  • 56218 02/06/01 15:24 Saha HAC Display - LVDT pointing readings
  • Monday

  • 56217 02/05/01 14:12 R. Michaels trigger temp. disabled
  • Saturday

    
    

    Thursday

  • 56214 02/01/01 00:12 Lingyan reason of decoder error: VDC threhold zero